# Structured Generative Modeling with The Thermodynamic Kolmogorov-Arnold Model

### v3

@@ A.1 Codebase
The study was implemented at https://github.com/exa-laboratories/T-KAM.
@@ A.2 Computational benchmarks
These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce GTX 1650) with Julia’s BenchmarkTools package, (JuliaCI, 2024). The index search operations in Eqs. 2313, & 70 were parallelized across 12 threads on a CPU (i7-9750H @ 2.60GHz).
Figure 8: Benchmarks for varying latent dimensions, $n_{z}$, while sampling from independent EBM priors with ITS and estimating posterior expectations with IS. The prior had an input size of $n_{z}$ and an output size of $2n_{z}+1$. The generator received inputs of size $2n_{z}+1$, with a hidden size of $4n_{z}+2$ and an output size of $28\times 28$. All other hyperparameters were kept consistent with Tab. 4. They reflect the performance of a reverse autodifferentiation of the MLE criterion, Eq. 16.
Figure 9: Benchmarks for varying amounts of ULA iterations, $N_{prior}$, to sample from a deep EBM prior (Tab. 8). Posterior expectations were estimated with an Monte Carlo estimator of the MLE criterion, Eq. 16. All other hyperparameters were kept consistent with Tab. 6. They reflect the computational cost of generating samples from the prior with Eq. 33. The benchmarks for ULA appear relatively constant across different step counts due to Julia’s loop optimizations and efficient memory management. Nevertheless, the reference ITS sampler remains significantly faster in practice. Note: Since the reference ITS sampling executed very rapidly, its relatively high proportion of time attributed to garbage collection is not unexpected.
Figure 10: Benchmarks for varying amounts of temperatures, $N_{t}$, while using a mixture of EBM priors (Tab. 7) with ITS sampling. Posterior expectations were estimated with SE and population-based ULA. All other hyperparameters were kept consistent with Tab. 6. They reflect the computational cost of performing a single reverse autodifferentiation of the SE criterion, Eq. 31. Note: Since the reference MLE criterion is executed very rapidly, its relatively high proportion of time attributed to garbage collection is not unexpected.
@@ A.3.1 Cyclic annealing
Instead, we schedule $p$ cyclically across all parameter updates of T-KAM’s training procedure, similar to the previous work of Fu et al. (2019). The initial and final $p$’s and the number of cycles are treated as tunable hyperparameters:
$\delta$ $=\text{linspace}\left(0,\ 2\pi(N_{\text{cycles}}+0.5);\ \text{len}=N_{\text{updates}}\right)$
$p(i)$ $=p_{\text{init}}+\frac{p_{\text{end}}-p_{\text{init}}}{2}\left(1-\cos\delta\right)$ (37)
where $N_{\text{updates}}$ denotes the total number of parameter updates, $i$ indexes the current training iteration, $N_{\text{cycles}}$ is the number of cycles, and $\delta$ is a vector of length $N_{\text{updates}}$.
@@ A.3.2 Choosing hyperparameters
${N_{t}}>1$ Specifies the number of power posteriors, enabling a smooth transition between the prior and posterior. Increasing this improves exploration but also increases cost due to the need for more MALA iterations. Typically, this is limited and therefore guided by hardware availability, such as $N_{t}=5$.
$N_{\text{local}}>1$ Determines the number of ULA or autoMALA iterations per power posterior. Although increasing this value improves the quality of samples at each stage, it also raises the computational cost, much like $N_{t}$. It can be kept at a moderate value, such as $N_{\text{local}}=20$, (based on the recommendations of Nijkamp et al. (2019b; a) regarding short-run Markov chain Monte Carlo).
@@
$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}$ Controls the evolution of tempering during training. If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$. For data with finer details, increase both values to allow for better focus on capturing intricate features. Recommended values are $p_{\text{init}}=4,$ and $p_{\text{end}}=1$, since our experience suggests that power posteriors change rapidly closer to the prior.
@@ A.5 Hyperparameters and training details
layer_norm (-) true
@@
generator_variance $\sigma_{\epsilon}$ 1.0
@@
num_grid_updating_samples (-) 50
@@
Training examples $N_{\text{train}}$ 20,000
@@
layer_norm (-) true
@@
generator_variance $\sigma_{\epsilon}$ 0.1
@@
batch_size $N_{x}$ 50
importance_sample_size $N_{s}$ 50
@@
Number of epochs $N_{\text{epochs}}$ 600
@@
spline_activation (-) Silu
@@
spline_degree $d$ 3
grid_size $N_{\text{grid}}$ 0, (Chebyshev requires no grid)
@@
grid_range $z_{\text{min}},z_{\text{max}}] [$1.2,1.2][-1.2,1.2]
@@
Initial basis param. $\tau_{0}$ 1.01
@@
GaussQuad_nodes $N_{\text{quad}}$ 100
Use mixture prior (-) variable
@@
All other KAN hparams N/A
generator_variance $\sigma_{\epsilon}$ 0.3
@@
update_prior_grid (-) false
@@
Training examples $N_{\text{train}}$ 40,000
@@
Number of epochs $N_{\text{epochs}}$ 20
@@
PRIOR_LANGEVIN
iters $N_{\text{local}}$ 60
step_size $\eta$ 0.16
@@
iters $N_{\text{local}}$ SVHN: 20
CIFAR10: 40
step_size $\eta$ 0.01
@@
iters $N_{\text{local}}$ 20
num_temps $N_{t}$ 5
Initial exponent in power law schedule $p_{\text{start}}$ 4
Final exponent in power law schedule $p_{\text{end}}$ 1
num_cycles $N_{\text{cycles}}$ 0
@@
A.5.1 SVHN & CIFAR-10 Architectures
Table 7: Dense Chebyshev KAN architecture for mixture prior model
@@ A.5.1 SVHN & CIFAR-10 Architectures
Input Latent vector $\bm{z}$ $p=100$
100 Dense($100$) $q=201$
@@
Table 8: Dense Chebyshev KAN architecture for deep prior model
Layer Operation Output Dim.
Input Latent vector $\bm{z}$ $p=100$
100 Dense($100$) $200$
200 Dense($200$) $200$
200 Dense($200$) $q=1$
sizes=[100,200,200,1]
Table 9: CNN Generator Architecture for SVHN
@@ 3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
Config: kernel=4, strides=1/2/2/2, paddings=0/1/1/1, activation=LeakyReLU, batchnorm=False
Table 10: CNN Generator Architecture for CIFAR-10
Layer Operation Output Shape Stride
Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
1 $\times8$ convT($1024$), LeakyReLU $8\times 8\times 1024$ 1
2 $\times4$ convT($512$), LeakyReLU $16\times 16\times 512$ 2
3 $\times4$ convT($256$), LeakyReLU $32\times 32\times 256$ 2
4 $\times3$ convT(3), Sigmoid $32\times 32\times 3$ 1
Config: kernel=8/4/4/3, strides=1/2/2/1, paddings=0/1/1/1, activation=LeakyReLU, batchnorm=False
@@ A.6 Sampling theory
This section provides more detail regarding T-KAM’s sampling procedures.
@@ A.6.3 Metropolis-adjusted Langevin algorithm (MALA)
A.7 $20\times 20$ image grids
@@ A.7.1 MNIST
Figure 13: Generated MNIST, (Deng, 2012), after 2,000 parameter updates using MLE / IS adhering to KART’s structure. Uniform, lognormal, and Gaussian priors are contrasted using Radial Basis Functions, (Li, 2024). Lognormal is inverted for clarity.
Figure 14: Generated MNIST, (Deng, 2012), after 2,000 parameter updates using MLE / IS adhering to KART’s structure. Uniform, lognormal, and Gaussian priors are contrasted using Fourier Bases, (Xu et al., 2024). Lognormal is inverted for clarity.
@@ A.7.2 FMNIST
Figure 15: Generated FMNIST, (Xiao et al., 2017), after 2,000 parameter updates using MLE / IS adhering to KART’s structure. Uniform, lognormal, and Gaussian priors are contrasted using Radial Basis Functions, (Li, 2024). Lognormal is inverted for clarity.
Figure 16: Generated FMNIST, (Xiao et al., 2017), after 2,000 parameter updates using MLE / IS adhering to KART’s structure. Uniform, lognormal, and Gaussian priors are contrasted using Fourier Bases, (Xu et al., 2024). Lognormal is inverted for clarity.
@@ A.7.3 Darcy flow
Figure 17: Generated 2D Darcy flow pressures, (Li et al., 2021), after 12,000 parameter updates using MLE / IS while adhering to KART. Radial Basis Functions, (RBFs) (Li, 2024), are contrasted against Fourier bases, (FFTs) (Xu et al., 2024), using Gaussian latent priors. RBF is colored differently for clarity.
Figure 18: Generated 2D Darcy flow pressures, (Li et al., 2021), after 12,000 parameter updates using MLE / IS while adhering to KART. Radial Basis Functions, (RBFs) (Li, 2024), are contrasted against Fourier bases, (FFTs) (Xu et al., 2024), using lognormal latent priors. RBF is colored differently for clarity.
Figure 19: Generated 2D Darcy flow pressures, (Li et al., 2021), after 12,000 parameter updates using MLE / IS while adhering to KART. Radial Basis Functions, (RBFs) (Li, 2024), are contrasted against Fourier bases, (FFTs) (Xu et al., 2024), using uniform latent priors. RBF is colored differently for clarity.
@@ A.7.4 SVHN
(a) SVHN with mixture prior and MLE
(b) SVHN with mixture prior and SE
Figure 20: Generated samples from T-KAM trained on SVHN using a mixture Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ITS prior and ULA posterior sampling.
(a) SVHN with deep prior and MLE
(b) SVHN with deep prior and SE
Figure 21: Generated samples from T-KAM trained on SVHN using a deep Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ULA prior and ULA posterior sampling.
A.7.5 CIFAR-10
(a) CIFAR-10 with mixture prior and MLE
(b) CIFAR-10 with mixture prior and SE
Figure 22: Generated samples from T-KAM trained on CIFAR-10 using a mixture Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ITS prior and ULA posterior sampling.
(a) CIFAR-10 with deep prior and MLE
(b) CIFAR-10 with deep prior and SE
Figure 23: Samples from T-KAM trained on CIFAR-10 using a deep Chebyshev KAN prior, (SS et al., 2024), after 8,000 parameter updates with ULA prior and ULA posterior sampling.
A.8 Top 3 extracted priors
@@ A.8.1 MNIST RBF
Figure 24: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 25: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 26: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.2 MNIST FFT
Figure 27: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 28: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 29: Three components from T-KAM’s prior after training on MNIST for 2,000 parameter updates with FFT bases and uniform constrained priors.
@@ A.8.3 FMNIST RBF
Figure 30: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 31: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 32: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.4 FMNIST FFT
Figure 33: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 34: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 35: Three components from T-KAM’s prior after training on FMNIST for 2,000 parameter updates with FFT bases and uniform constrained priors.
@@ A.8.5 Darcy RBF
Figure 36: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 37: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 38: Three components from T-KAM’s prior after training on Darcy flow pressures for 12,000 parameter updates with RBF bases and uniform constrained priors.

### v4

@@ A.1 Codebase
The study was implemented at https://github.com/thezettascale/T-KAM.
@@ A.2 Computational benchmarks
These benchmarks were collected for FP32 on a GPU (NVIDIA GeForce RTX 4060) with Julia’s BenchmarkTools package, (JuliaCI, 2024). The index search operations in Eqs. 2313, & 69 were parallelized across 20 threads on a CPU (i9-13900H @ 5.40GHz).
Figure 7: Benchmarks for varying latent dimensions, $n_{z}$, while sampling from independent EBM priors with ITS and estimating posterior expectations with IS. The prior had an input size of $n_{z}$ and an output size of $2n_{z}+1$. The generator received inputs of size $2n_{z}+1$, with a hidden size of $4n_{z}+2$ and an output size of $28\times 28$. All other hyperparameters were consistent with Tab. 4 (RBF bases). They reflect the performance of a reverse autodifferentiation of the MLE criterion, Eq. 16.
Figure 8: Benchmarks for varying amounts of temperatures, $N_{t}$, using a mixture of EBM priors (Tab. 7) with ITS sampling. Posterior expectations were estimated with SE and population-based ULA. All other hyperparameters were kept consistent with Tab. 6 (Celeb-A). They reflect the computational cost of performing a single reverse autodifferentiation of the SE criterion, Eq. 31.
@@ A.3.1 Cyclic annealing
Instead, we schedule $p$ sinusoidally across all training parameter updates, starting with a higher value of $p$ and ending with a lower value, treated as tunable hyperparameters.
Figure 11: Cosine $p$ schedule, decreasing from 10 to 1
@@ A.3.2 Choosing hyperparameters
${N_{t}}>1$ Specifies the number of power posteriors, enabling a smooth transition between the prior and posterior. Increasing this improves exploration but also increases cost due to the need for more MALA iterations. Typically, this is memory-bound and guided by hardware availability, such as $N_{t}=5$.
$N_{\text{local}}>1$ Determines the number of ULA or autoMALA iterations per power posterior. Although increasing this value improves the quality of samples at each stage, it also raises the computational cost, much like $N_{t}$. It can be kept at a small value, such as $N_{\text{local}}=40$, (based on the recommendations of Nijkamp et al. (2019b; a) regarding short-run Markov chain Monte Carlo).
@@
$p_{\text{init}}>1,$ $p_{\text{end}}\leq 1,$ $N_{\text{cycles}}>0$ Controls the evolution of tempering during training. If the data is expected to have smooth relationships, reduce both $p_{\text{init}}$ and $p_{\text{end}}$. For data with finer details, increase both values to allow for better focus on capturing intricate features. Recommended values are $p_{\text{init}}=10,$ and $p_{\text{end}}=3$, since our experience suggests that power posteriors change rapidly closer to the prior. The number of cycles, $N_{\text{cycles}}$, controls the periodicity of cosine scheduling similar to cyclic annealing (Fu et al., 2019).
@@ A.5 Hyperparameters and training details
layer_norm (-) false
@@
generator_noise $\sigma_{\epsilon}$ 0.1
generator_variance $\sigma_{\text{llhood}}$ 0.1
@@
num_grid_updating_samples (-) 100
@@
Training examples $N_{\text{train}}$ 50,000
@@
layer_norm (-) false
@@
generator_noise $\sigma_{\epsilon}$ 0.1
generator_variance $\sigma_{\text{llhood}}$ 0.1
@@
batch_size $N_{x}$ 100
importance_sample_size $N_{s}$ 100
@@
Number of epochs $N_{\text{epochs}}$ 1,000
@@
spline_activation (-) None
@@
spline_degree $d$ 5
grid_size $N_{\text{grid}}$ 1
@@
grid_range $z_{\text{min}},z_{\text{max}}] [$1,1][-1,1]
@@
Initial basis param. $\tau_{0}$ 1.001
@@
GaussQuad_nodes $N_{\text{quad}}$ 200
Use mixture prior (-) true
@@
All other KAN hparams N/A, an MLP CNN was used
generator_noise $\sigma_{\epsilon}$ 0.01
generator_variance $\sigma_{\epsilon}$ 0.25
@@
update_prior_grid (-) true
@@
Training examples $N_{\text{train}}$ 20,000
@@
Number of epochs $N_{\text{epochs}}$ 40
@@
iters $N_{\text{local}}$ 40
step_size $\eta$ 0.002
@@
iters $N_{\text{local}}$ (see posterior langevin)
num_temps $N_{t}$ 10
Initial exponent in power law schedule $p_{\text{start}}$ 10
Final exponent in power law schedule $p_{\text{end}}$ 3
Number of cosine cycles $N_{\text{cycles}}$ 0
@@
A.5.1 SVHN & CelebA Architectures
Table 7: The dense Chebyshev KAN network used as the mixture prior model
@@ A.5.1 SVHN & CelebA Architectures
Input Sample $\bm{z}$ $p=20$
Energy functions Dense(2020) $q=41$
@@
Table 8: CNN Generator Architecture for SVHN ($32\times 32$)
Layer Operation Output Shape Stride
Input Latent vector $\bm{z}$ reshaped $1\times 1\times q$
1 $\times4$ convT($256$), LeakyReLU $4\times 4\times 256$ 1
2 $\times4$ convT($128$), LeakyReLU $8\times 8\times 128$ 2
3 $\times4$ convT($64$), LeakyReLU $16\times 16\times 64$ 2
4 $\times4$ convT(3), Sigmoid $32\times 32\times 3$ 2
Config: kernel=4,4,4,4 strides=1/2/2/2, paddings=0/1/1/1, activation=LeakyReLU, batchnorm=False
Table 9: CNN Generator Architecture for CelebA ($64\times 64$)
@@ 3 $\times4$ convT($128$), LeakyReLU $16\times 16\times 128$ 2
4 $\times4$ convT(64), LeakyReLU $32\times 32\times 64$ 2
5 $\times4$ convT(3), Sigmoid $64\times 64\times 3$ 2
Config: kernel=4,4,4,4,4 strides=1/2/2/2/2, paddings=0/1/1/1/1, activation=LeakyReLU, batchnorm=False
@@ A.6 Sampling theory
This section provides more theoretical detail regarding T-KAM’s sampling procedures.
@@ A.6.3 Metropolis-adjusted Langevin algorithm (MALA)
A.7 Extra image samples
@@ A.7.1 MNIST
(a) Real samples
(b) No ref. prior, RBF basis
(c) No ref. prior, FFT basis
(d) Gaussian ref. prior, RBF basis
(e) Gaussian ref. prior, FFT basis
(f) Uniform ref. prior, RBF basis
(g) Uniform ref. prior, $F$ basis
(h) Lognormal prior, RBF basis
(i) Lognormal ref. prior, FFT basis
Figure 12: Generated MNIST examples using different bases and exponentially-tilted priors (Sec. 3.1.1).
@@ A.7.2 FMNIST
(a) Real samples
(b) No ref. prior, RBF basis
(c) No ref. prior, FFT basis
(d) Gaussian ref. prior, RBF basis
(e) Gaussian ref. prior, FFT basis
(f) Uniform ref. prior, RBF basis
(g) Uniform ref. prior, $F$ basis
(h) Lognormal prior, RBF basis
(i) Lognormal ref. prior, FFT basis
Figure 13: Generated FMNIST examples using different bases and exponentially-tilted priors (Sec. 3.1.1).
@@ A.7.3 Darcy flow
(a) Real samples
(b) No ref. prior, RBF basis
(c) No ref. prior, FFT basis
(d) Gaussian ref. prior, RBF basis
(e) Gaussian ref. prior, FFT basis
(f) Uniform ref. prior, RBF basis
(g) Uniform ref. prior, $F$ basis
(h) Lognormal prior, RBF basis
(i) Lognormal ref. prior, FFT basis
Figure 14: Generated Darcy flow examples using different bases and exponentially-tilted priors (Sec. 3.1.2).
@@ A.7.4 SVHN
(a) Maximum Likelihood
(b) Steppingstone estimator
Figure 15: Generated SVHN examples by T-KAM trained with MLE and SE.
A.7.5 CelebA
(a) Maximum Likelihood
(b) Steppingstone estimator
Figure 16: Generated CelebA examples by T-KAM trained with MLE and SE.
A.8 Trained latent priors
@@ A.8.1 MNIST RBF
Figure 17: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 18: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 19: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.2 MNIST FFT
Figure 20: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 21: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 22: Three components from T-KAM’s prior after training on MNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
@@ A.8.3 FMNIST RBF
Figure 23: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 24: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 25: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with RBF bases and uniform constrained priors.
@@ A.8.4 FMNIST FFT
Figure 26: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and Gaussian constrained priors.
@@
Figure 27: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and lognormal constrained priors.
@@
Figure 28: Three components from T-KAM’s prior after training on FMNIST for 5,000 parameter updates with FFT bases and uniform constrained priors.
@@ A.8.5 Darcy RBF
Figure 29: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and Gaussian constrained priors.
@@
Figure 30: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and lognormal constrained priors.
@@
Figure 31: Three components from T-KAM’s prior after training on Darcy flow pressures for 10,000 parameter updates with RBF bases and uniform constrained priors.